# Operation of the Trigger System for the ICARUS Detector at Fermilab

### v1

@@ 3.3 Analysis of the collected data
  • Stopping muons as recognized by the increased $\text{d}E/\text{d}x$ local energy depositions at the end of the tracks, to characterize the trigger as function of the energy. They are uniformly distributed across the detector and are characterized by an energy spectrum in the same range as BNB charged-current muon neutrino interactions (see figures 1 and 12), with roughly the same event spatial extension in the detector.
@@
Figure 11: Example of event in the ICARUS West module from Run2 minimum bias data. The TPC is read out for $1.64$ms, and the reconstructed tracks are matched with the CRT, which is read out for $3$ms. Some of the tracks are clipped due to the TPC readout opening (closing) after (before) the particle enters the active volume of the detector. Tracks selected with the vertical and stopping selections are highlighted with green and red shades respectively.
Examples of selected through-going and stopping muons are visualized in figure 11, with the corresponding $\text{d}E/\text{d}x$ local energy depositions measured with granularity at millimeter level. While the deposited energy for through-going muons is measured from the charge collected on the TPC wires with a $\approx 5\%$ resolution [16], the energy for stopping muons is more precisely determined from their reconstructed residual range. For this purpose, the ionization charge was calibrated by accounting for the free drift electron lifetime in LAr ($\approx 5$ms in the analyzed data sample) [14], and electron-ion recombination [16].
@@
Figure 12 (left) shows the distribution of the $\text{d}E/\text{d}x$ with respect to the residual range for the selected stopping muons, along with the Bethe-Bloch prediction for the mean $\text{d}E/\text{d}x$ value. The typical ascent at the end of each track is evident, highlighting the purity of the selected sample. The corresponding energy spectrum, measured from the reconstructed range, is reported in figure 12 (right).
@@
Muon tracks longer than $20$cm were matched to CRT signals with $d_{\text{CRT}}<30\,\text{cm}$, within the PMT readout window. A one-to-one match was required between each CRT hit and track, to reduce the contamination by tracks and showers reconstructed as multiple collinear tracks that are matched to the same CRT hit, which degrades the reconstruction of the tracks length and deposited energy. In this way, muons were selected independently of their orientation (“loose selection”). In addition, vertical muon tracks were selected entering the detector at a $<20^{\circ}$ angle with respect to the vertical direction, and matched only to Top CRT signals. These tracks were allowed to have the start (end) point up to within $15$cm from the top (bottom) TPC wall to accommodate for distortions from electric field non-uniformity, and were required to have a reconstructed length in agreement within 3% with the start-to-end distance. The reconstructed length for the selected vertical muons is shown in figure 13.
@@ 3.4 Trigger efficiency measurements
The majority-5 (Mj=5, see section 2.3) in-spill trigger efficiency on stopping muons is reported in figure 14 as a function of the muon energy determined from the residual range ($>70$MeV). The efficiency roughly saturates at $400$MeV both in Run1 and Run2, with the plateau largely covering the spectrum of charged-current neutrino interactions from the BNB and the higher-energy NuMI beams (see figure 1). The introduction of the two overlapped windows in Run2 results in a $\approx 10\%$ improvement below $300$MeV compared to Run1, where the detector was logically divided in only three regions (see figure 7). The trigger efficiency in the East module shows a reduction by up to 3% compared to the West module(figure 14), due to the presence of 11 and 3 dead PMT channels in the two modules respectively.
Figure 14: Mj=5 trigger efficiency with stopping muons as function of the energy from the residual muon range in Run1 and Run2 (left), and separately in the two West and East modules for Run2 (right). The average simulated deposited energy for BNB $\nu_{\mu}$ charged-current interactions is indicated for reference.
[…]
@@
(section 3.1). The trigger emulation software looked for PMT signals in a $300$ns window around the time of the muon as assigned with the CRT. A CRT-TPC mismatch would lead to a wrong evaluation of the track time. This source of systematics was removed by requiring a loose spatial correspondence of a minimal light signal on few PMTs with the track charge deposition in the TPC. In the case of cathode crossing tracks, the time assigned from the CRT was requested to match the one assigned from the TPC (see section 3.3).
@@
The stopping muon selection was based on the identification of the Bragg peak by the measured ionization signal, which would be affected by an incomplete reconstruction of the track. A dedicated Monte Carlo study indicated that most tracks were correctly reconstructed, with a resolution on the length better than $1\%$. The selection mitigated the impact of broken tracks (see section 3.3), which would lead to an underestimation of the energy from the residual range and to an overestimation of the trigger efficiency. The residual lower-bound systematic uncertainty on the efficiency amounts to 11% below $200$MeV, 1% between 200 and $300$MeV, and is negligible everywhere else.
@@
The software emulation of the trigger response, applied to a dedicated minimum bias data sample in which additional trigger hardware information was added, showed $\approx 5\%$ of events with an emulated trigger not confirmed by the hardware. The tracks generating such emulated triggers were typically located at the corners of the detector and had low deposited energy, yielding weak and short light signals which might have been missed by the misconfigured hardware (see section 3.2). In Run1, the efficiency shows a reduction of $11\%$ below $200$MeV, $5.5\%$ between 200 and $300$MeV, $4.7\%$ between 300 and $400$MeV and $<1\%$ at higher energies. In Run2, the decrease is of $14\%$ below $200$MeV, $8\%$ between 200 and $300$MeV, $3.7\%$ between 300 and $400$MeV and $<1\%$ above $400$MeV. The trigger efficiency has been corrected accordingly, with a conservative systematic uncertainty amounting to half of the correction.
$<200$MeV $[200,300]$MeV $[300,400]$MeV $>400$MeV
Statistical 6% 2% $<1\%$ $<1\%$
Systematic 11% 4% $2\%$ $<1\%$
@@
The spatial uniformity of the trigger response was probed with the selected vertical muons (see section 3.3) in the $[0.6,1.0]$GeV deposited energy range. The event recognition efficiency with the emulated Mj=5 trigger is roughly saturated (figure 15) in both Run1 and Run2, regardless of the muon position.
@@
The corresponding trigger efficiency for collecting light from out-of-spill cosmic rays satisfying Mj=10 (Mj=9) in Run1 (Run2) is shown in figure 16, using a muon sample selected loosely, independently of the particle orientation (see section 3.3). Compared to Run1, the Run2 out-of-spill trigger efficiency is improved also in uniformity, by including the two additional overlapping windows into the trigger logic (section 2.3).
@@ 4 Conclusions
The trigger efficiency for a majority-5 of fired PMT pairs is found to be almost full for stopping muons above $300$MeV in both modules, with remarkable spatial uniformity, and it reaches up to $>80\%$ between 200 and $300$MeV. Analogously, the efficiency of the stricter, majority-9 trigger, used in Run2 to identify cosmic rays crossing the detector during the $1$ms TPC drift time, is found to be $>80\%$ for $>400$MeV energy deposition.

### v2

@@ 3.3 Analysis of the collected data
  • Stopping muons as recognized by the increased $\text{d}E/\text{d}x$ local energy depositions at the end of the tracks, to characterize the trigger as function of the energy. They are uniformly distributed across the detector and are characterized by an energy spectrum in the same range as BNB charged-current muon neutrino interactions (see figures˜1 and 12), with roughly the same event spatial extension in the detector.
@@
Figure 11: Example of event in the ICARUS West module from Run2 minimum bias data. The TPC is read out for $1.64$ ms, and the reconstructed tracks are matched with the CRT, which is read out for 33 ms. Some of the tracks are clipped due to the TPC readout opening (closing) after (before) the particle enters the active volume of the detector. Tracks selected with the vertical and stopping selections are highlighted with green and red shades respectively.
Examples of selected through-going and stopping muons are visualized in figure˜11, with the corresponding $\text{d}E/\text{d}x$ local energy depositions measured with granularity at millimeter level. While the deposited energy for through-going muons is measured from the charge collected on the TPC wires with a $\approx 5\%$ resolution [17], the energy for stopping muons is more precisely determined from their reconstructed residual range. For this purpose, the ionization charge was calibrated by accounting for the free drift electron lifetime in LAr ($\approx 5$ ms in the analyzed data sample) [15], and electron-ion recombination [17].
@@
Figure˜12 (left) shows the distribution of the $\text{d}E/\text{d}x$ with respect to the residual range for the selected stopping muons, along with the Bethe-Bloch prediction for the mean $\text{d}E/\text{d}x$ value. The typical ascent at the end of each track is evident, highlighting the purity of the selected sample. The corresponding energy spectrum, measured from the reconstructed range, is reported in figure˜12 (right).
@@
Muon tracks longer than $20$ cm were matched to CRT signals with $d_{\text{CRT}}<30\,\text{cm}$, within the PMT readout window. A one-to-one match was required between each CRT hit and track, to reduce the contamination by tracks and showers reconstructed as multiple collinear tracks that are matched to the same CRT hit, which degrades the reconstruction of the tracks length and deposited energy. In this way, muons were selected independently of their orientation (“loose selection”). In addition, vertical muon tracks were selected entering the detector at a $20^{\circ}$ angle with respect to the vertical direction, and matched only to Top CRT signals. These tracks were allowed to have the start (end) point up to within $15$ cm from the top (bottom) TPC wall to accommodate for distortions from electric field non-uniformity, and were required to have a reconstructed length in agreement within 3% with the start-to-end distance. The reconstructed length for the selected vertical muons is shown in figure˜13.
@@ 3.4 Trigger efficiency measurements
The majority-5 (Mj=5, see section˜2.3) in-spill trigger efficiency on stopping muons is reported in figure˜14 as a function of the muon energy determined from the residual range (>70>70 MeV). The efficiency roughly saturates at $400$ MeV both in Run1 and Run2, with the plateau largely covering the spectrum of charged-current neutrino interactions from the BNB and the higher-energy NuMI beams (see figure˜1). The introduction of the two overlapped windows in Run2 results in a $\approx 10\%$ improvement below $300$ MeV compared to Run1, where the detector was logically divided in only three regions (see figure˜7). The trigger efficiency in the East module shows a reduction by up to 3% compared to the West module (figure˜14), due to the presence of 11 and 3 dead PMT channels in the two modules respectively.
Figure 14: Mj=5 trigger efficiency with stopping muons as function of the energy from the residual muon range in Run1 and Run2 (left), and separately in the two West and East modules for Run2 (right). The error bars include both systematic and statistic uncertainties added in quadrature. The average simulated deposited energy for BNB $\nu_{\mu}$ charged-current interactions is indicated for reference.
Uncertainties on the measured trigger efficiency were evaluated with the same approach in both Run1 and Run2, where they are of similar magnitude. The statistical term ranges up to $\approx 6\%$ below $200$ MeV energy of the stopping muons, $\approx 2\%$ between $200\,\text{MeV}$ and $300$ MeV, and $1\%$ at higher energy. Associated systematics come from the track reconstruction and selection, and from the emulation of the trigger PMT-majority logic using the recorded PMT signals:
@@
(section˜3.1). The trigger emulation software looks for PMT signals in a $300$ ns window around the time of the muon as assigned with the CRT. A CRT-TPC mismatch would lead to a wrong evaluation of the track time. This source of systematics was removed by requiring a loose spatial correspondence of a minimal light signal on few PMTs with the track charge deposition in the TPC. In the case of cathode crossing tracks, the time assigned from the CRT was requested to match the one assigned from the TPC (see section˜3.3).
@@
The stopping muon selection, based on the identification of the Bragg peak by the measured ionization signal (see section˜3.3), allowed to correctly reconstruct the large majority of tracks, with a resolution on the length better than $1\%$ as indicated by a dedicated Monte Carlo study. In the few remaining cases where the reconstruction of the tracks is incomplete, tracks may be broken and wrongly recognized as stopping muons, e.g., due to the presence of kinks and delta rays. In this case, the track length and associated energy would be underestimated, leading to an overestimation of the trigger efficiency at low energies. The impact of track mis-reconstruction was evaluated via Monte Carlo and included in the measurement of the trigger efficiency as a lower-bound systematic uncertainty of 11% below $200$ MeV, 1% between 200 and $300$ MeV, and negligible everywhere else.
@@
The software emulation of the trigger response, applied to a dedicated minimum bias data sample in which additional trigger hardware information was added, showed $\approx 5\%$ of events with an emulated trigger not confirmed by the hardware. The tracks generating such emulated triggers were typically located at the corners of the detector and had low deposited energy, yielding weak and short light signals which might have been missed by the misconfigured hardware (see section˜3.2). In Run1, the efficiency shows a reduction of $11\%$ below $200$ MeV, $5.5\%$ between 200 and $300$ MeV, $4.7\%$ between 300 and $400$ MeV and $1\%$ at higher energies. In Run2, the decrease is of $14\%$ below $200$ MeV, $8\%$ between 200 and $300$ MeV, $3.7\%$ between 300 and $400$ MeV and $1\%$ above $400$ MeV. The trigger efficiency has been corrected accordingly, with a conservative systematic uncertainty amounting to half of the correction.
Overall, as shown in table˜2 the combined systematic uncertainties amount to $\approx+7/-13\%$ below $200$ MeV, $\approx 4\%$ between 200 and $300$ MeV, $\approx 2\%$ between 300 and $400$ MeV, and are almost negligible at higher energies, where tracks are better recognized and measured.
<200<200 MeV $[200,300]$ MeV $[300,400]$ MeV >400>400 MeV
Statistical $6\%$ $2\%$ $\% <1<1\%$
Systematic $7\%/-13\%$ $\pm 4\%$ $\pm 2\%$ $\pm 1\%$
@@
The spatial uniformity of the trigger response was probed with the selected vertical muons (see section˜3.3) in the $[0.6,1.0]$ GeV deposited energy range. The event recognition efficiency with the emulated Mj=5 trigger is roughly saturated (figure˜15) in both Run1 and Run2, regardless of the muon position.
@@
The corresponding trigger efficiency for collecting light from out-of-spill cosmic rays satisfying Mj=10 (Mj=9) in Run1 (Run2) is shown in figure˜16, using a muon sample selected loosely, independently of the particle orientation (see section˜3.3). Compared to Run1, the Run2 out-of-spill trigger efficiency is improved also in uniformity, by including the two additional overlapping windows into the trigger logic (section˜2.3).
@@ 4 Conclusions
The trigger efficiency for a majority-5 of fired PMT pairs is found to be almost full for stopping muons above $300$ MeV in both modules, with remarkable spatial uniformity, and it reaches up to $80\%$ between 200 and $300$ MeV. Analogously, the efficiency of the stricter, majority-9 trigger, used in Run2 to identify cosmic rays crossing the detector during the $1$ ms TPC drift time, is found to be $80\%$ for >400>400 MeV energy deposition.